OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (7, -2) and has a slope of -3. y - 2 = -3(x + 7) y + 7 = -3(x - 2) y + 2 = -3(x - 7) y - 7 = -3(x + 2)

OpenStudy (compassionate):

So we have y - y1 = m(x - x1) So we are given (7, -2) and -3 y - (-2) = -3(x - 7) y + 2 = -3(x - 7) I hope this helped! :)